# A Noisy Approach to Intrinsically Mixed-State Topological Order

### v2

@@ III.1 $\mathbb{Z}_{2}^{(0)}$ and $\mathbb{Z}_{2}^{(1)}$ TSSC from $\mathbb{Z}_{2}$ Toric code
As the paradigmatic example of a topological stabilizer code, the stability of the $\mathbb{Z}_{2}$ Toric code to decoherence has been extensively investigated [14, 17, 19]. We revisit this problem here in light of our interpretation of the decohered state as a TSSC. Consider a system of qubits placed on the edges of a square lattice with periodic boundary conditions, with the Hamiltonian given by
@@
Recall that since $A_{s}\rho=\rho A_{s}=\rho$ and $B_{p}\rho=\rho B_{p}=\rho$, ground states of Eq. (8) can be interpreted as closed loop condensates of the $e$, $m$, and $f$ anyons. After maximal decoherence, we instead only have $B_{p}\rho_{e}=\rho_{e}B_{p}=\rho_{e}$ and $A_{s}\rho_{e}A_{s}=\rho_{e}$. Physically, the $e$-noise has the effect of “freezing” the $m$-loops (and hence also the $f$ loops) into a classical ensemble, while leaving the “quantum” condensate of $e$-loops untouched. More precisely, $e$-errors break the strong 1-form magnetic symmetry of the original pure state down to a weak 1-form symmetry, while leaving the strong 1-form electric symmetry intact. We will later provide a general discussion of the role 1-form symmetries play in characterizing generic imTOs (see Sec. V).
@@
Similar considerations hold for $\rho_{f}$. We have that $A_{v}B_{v-\mathbf{y}}\rho_{f}=\rho_{f}A_{v}B_{v-\mathbf{y}}=\rho_{f}$, where $A_{v}B_{v-\mathbf{y}}$ generates a closed $f$ loop and $v-\mathbf{y}$ denotes the plaquette to the south-east of vertex $v$, while we only have that $A_{v}\rho_{f}A_{v}=\rho_{f}$. Thus, $f$-noise freezes both the $e$ and $m$ loops but leaves the $f$ loops untouched, such that $\rho_{f}$ describes a quantum condensate of fermionic excitations (stated otherwise, $\rho_{f}$ retains a strong 1-form symmetry): this is not expected to occur in the ground state of a gapped, local 2D Hamiltonian. Remarkably, despite starting with a bosonic topological order, decoherence has resulted in a state in which the sole deconfined quantum excitation is a fermion. In a rough sense, local decoherence allows one to “peel off” half of the original state.
@@
which yields the stabilizer group $\mathcal{S}_{f}=\langle A_{v}B_{v-\mathbf{y}},W_{x,y}^{f}\rangle$. This precisely describes a topological subsystem code describing the $\mathbb{Z}_{2}^{(1)}$ topological order which, again, is non-modular [36]. Like the previous case, $f$ is transparent (it braids trivially with itself) and hence the decohered state encodes no quantum memory, consistent with Ref. [27].
@@
It is readily apparent that the resulting TO in either case is that of a single $\mathbb{Z}_{2}$ Toric code, with the fusion group of the Abelian anyons given by $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. This can also be directly verified with the explicit forms of |$\ket{\rho_{e,f}}\rangle$ in the lattice model. In light of our above stabilizer analysis however, we note a key distinction between the mixed-states $\mathcal{C}_{e}$ and $\mathcal{C}_{f}$. Restricting attention to anyons with support on only the ket or bra space, we see that both orders support a single such anyon. For $\mathcal{C}_{e}$, this is the boson $e_{+}\sim e_{-}$, while for $\mathcal{C}_{f}$, this is the fermion $f_{+}\sim f_{-}$, where the equivalences are up to fusion with the condensed anyon. This is consistent with our observation in the stabilizer analysis that under $e$ and $f$ noise, the sole remaining deconfined quantum excitations are simply the original $e$ and $f$ excitations, respectively.
@@
We now show that this anyon condensation across the ket and bra spaces, at the level of the density matrix in the original Hilbert space, corresponds to gauging out an anyon. Recall that anyon condensation proceeds in two steps: to condense an anyon $a$, one first (i) projects out from the theory those anyons which braid non-trivially with $a$ (i.e. they become confined excitations) and then (ii) identifies those anyon types which differ by fusion with $a$. For instance, in the usual Toric code, condensing $e$ confines the $m$ and $f$ anyons while $e$ becomes identified with the vacuum: the resulting state has no remaining anyon excitations and is trivial. Gauging out an anyon however corresponds to only performing step (i) of this process. For instance, gauging out $e$ still confines $m$ and $f$, but leaves $e$ distinct from the vacuum, such that one is left with the anyon content $\{1,e\}$—precisely that of the $\mathbb{Z}_{2}^{(0)}$ non-modular TO realized via decoherence of $e$.
@@
Crucially, one can also gauge out anyons that cannot be condensed (non-bosonic anyons); analogously, one can decohere non-bosonic anyons $a$ since this corresponds to the conventional condensation of the bosonic pair $(a_{+}a_{-})$ in the doubled Hilbert space. For instance, one may gauge out $f$ from the $\mathbb{Z}_{2}$ Toric code to obtain the $\mathbb{Z}_{2}^{(1)}$ TO, precisely replicating the effect of $f$ errors. Surprisingly, as this simple example illustrates, locally correlated errors (which correspond to anyon condensation under the Choi map) provide a physical implementation of the gauging out procedure, which thus far remained a conceptual device for generating topological subsystem codes from topological stabilizer codes [28, 29, 30, 31, 36].
@@
Let us pause to recapitulate our observations in the context of the Toric code. We found that anyonic decoherence, previously shown to correspond to anyon condensation across the ket and bra spaces, implements the gauging out of anyons in the original Hilbert space, including those which are forbidden from condensing under purely unitary evolution. This process led to mixed-states supporting non-modular topological order, corresponding to topological subsystem codes, which is not believed to be allowed in the ground state of a locally gapped Hamiltonian in 2D. Moreover, the resulting mixed-state corresponds to a zero-correlation length state with strictly lower logical information than the parent state and thus represents a distinct mixed-state quantum phase of matter [19] i.e., it represents an intrinsically mixed topological phase of matter. Here, the stability of the maximally decohered state can be attributed to the fact that LPQCs can create at most area-law entanglement in finite time. For the specific case of the $\mathbb{Z}_{2}$ Toric code subject to $f$ errors, Ref. [27] numerically verified the robustness of the resulting imTO against finite noise channels that explicitly break the strong 1-form symmetry. However, we generically expect states with strong 1-form symmetries to remain robust up to some finite noise rate.
We now explore the above principles through examples that illustrate the general structures permitted in the decohered mixed-states. Specifically, we choose parent states with Abelian TO (that admit gapped boundaries) and identify appropriate error channels that lead to imTO with mixed quantum-classical memories as well as chiral anyon theories. Note that the presence of a classical memory is tied to non-local stabilizers in the decohered theory, which correspond to transparent anyons.
@@ III.2 $\mathbb{Z}_{4}^{(1)}$ TSSC from $\mathbb{Z}_{4}$ Toric code
$H_{\mathbb{Z}_{4}}=-\sum_{s}(A_{s}+A_{s}^{\dagger})-\sum_{p}(B_{p}+B_{p}^{\dagger}),$ (15)
with the star and plaquette operators defined in Fig. 2. The ground state manifold is determined by the constraints $A_{s}=B_{p}=1$, violations of which indicate the presence of electric $e$ and magnetic $m$ excitations, respectively. Explicitly, $Z_{i}$ applied to the ground state excites an $e$ and an $e^{-1}$ anyon at vertices connected by the edge $i$. Likwise, applying $X_{i}$ creates an $m$ and $m^{-1}$ on plaquettes separated by the edge $i$. These anyons satisfy $\mathbb{Z}_{4}\times\mathbb{Z}_{4}$ fusion rules $e^{4}=m^{4}=1$ and the braiding statistics between two composite objects $e^{a}m^{b}$ and $e^{c}m^{d}$ is given by $\mathcal{B}_{\theta}(ab,cd)=i^{ad+bc}$. On the torus, the non-contractible Wilson loops $W^{e}_{x,y}=\prod_{i\in\Gamma_{x,y}}Z_{i}$ and $W^{m}_{x,y}=\prod_{i\in\hat{\Gamma}_{x,y}}X_{i}$ serve as the logical operators and satisfy the algebra $W_{x/y}^{e}W_{y/x}^{m}=iW_{y/x}^{m}W_{x/y}^{e}$, such that the code space stores two $d=4$ qudits.
@@
Said more formally, $\hat{\mathcal{A}}$ errors break most of the strong 1-form symmetries of $\rho$ down to weak 1-form symmetries, while leaving the strong 1-form symmetries corresponding to $\mathcal{A}$ anyons intact. This is encoded in Eq. (6) and the fact that the stabilizer group $\mathcal{S}_{e^{-1}m}$ for the TSSC is generated by small loops for the $em$ anyon (which generates $\mathcal{A}$). Finally, since $e^{2}m^{2}$ is transparent in $\hat{\mathcal{A}}$, it remains transparent in $\mathcal{A}$ by definition. We thus obtain the same result as above: the set of decohered density matrices on the torus form the code space for a TSSC which describes a non-modular Abelian anyon theory $\mathcal{A}$, whose non-trivial anyons are two semions and a transparent boson. The mutual statistics of the semions result in this mixed-state encoding a logical qubit in its logical subsystem, and since it is again a zero-correlation length mixed-state (we are working at maximal dephasing), we expect it will be stable up to some finite noise threshold (for local noise).
It is instructive to once again consider the gauging out procedure from the perspective of the doubled Hilbert space. In this picture, decoherence of the anyons in $\hat{\mathcal{A}}$ corresponds to condensing $\{[1]_{+}[1]_{-},$ $[e^{-1}m]_{+}[e^{-1}m]_{-},[e^{2}m^{2}]_{+}[e^{2}m^{2}]_{-},[em^{3}]_{+}[em^{3}]_{-}\}$, which form a Lagrangian subgroup of the TO in the doubled space $\mathcal{C}\times\bar{\mathcal{C}}$, where $\mathcal{C}=\mathbb{Z}_{4}\times\mathbb{Z}_{4}$. Anyon condensation proceeds in the usual way: each anyon from $\hat{\mathcal{A}}$ is identified with the vacuum. Next, any excitation which braids non-trivially with any condensed anyon becomes confined and, of the remaining deconfined excitations, any that differ only up to fusion by anyons in $\hat{\mathcal{A}}$ are identified. A simple calculation shows that the resulting topological order is that of a $\mathbb{Z}_{4}$ gauge theory, with only the following anyons supported solely on the ket space: $[em]_{+}$, and $[e^{-1}m^{-1}]_{+}$, while $[e^{2}m^{2}]_{+}\sim[e^{2}m^{2}]_{-}$ can move freely between the ket and bra spaces, and is a transparent anyon. These are the deconfined excitations in the decohered theory (which correspond to strong 1-form symmetries of the resulting state), which unsurprisingly match those we obtained by working directly in the Hilbert space, where the anyons in $\hat{\mathcal{A}}$ were gauged out rather than condensed.
@@ III.3 Chiral Semion from Double Semion
As an instance of this mapping, let us take the doubled semion anyon theory as our parent TO. This theory can be realized as a Pauli stabilizer Hamiltonian [32] and its anyons form a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ group under fusion, with elements $\{1,s,\bar{s},s\bar{s}\}$. Here, $s$ is a semion (it has self-statistics $\theta(s)=i$), $\bar{s}$ is an anti-semion ($\theta(\bar{s})=-i$), and $s\bar{s}$ is a boson. Now, we subject a ground state of this system to an error channel that incoherently proliferates the semion $s$ i.e., $\hat{\mathcal{A}}=\{1,s\}$. This corresponds to gauging out $s$, which braids trivially with $\bar{s}$. The resulting anyon theory for the decohered mixed-state is given by $\mathcal{A}=\{1,\bar{s}\}$ i.e., it is the chiral (anti)-semion Abelian anyon theory. This zero-correlation length state represents an intrinsically mixed-state phase of matter as chiral UMTCs are widely held to be forbidden in the ground states of locally commuting Hamiltonians [46] and are not expected to be realized in fixed point wave-functions. Since $\bar{s}$ has non-trivial self-statistics, the decohered code space encodes exactly one logical qubit; this example represents the minimal model in which one obtains a chiral anyon theory with an encoded logical qubit.

### v3

@@ III.1 $\mathbb{Z}_{2}^{(0)}$ and $\mathbb{Z}_{2}^{(1)}$ TSSC from $\mathbb{Z}_{2}$ Toric code
As the paradigmatic example of a topological stabilizer code, the stability of the $\mathbb{Z}_{2}$ Toric code to decoherence has been extensively investigated [13, 15, 17, 19, 21]. We revisit this problem here in light of our interpretation of the decohered state as a TSSC. Consider a system of qubits placed on the edges of a square lattice with periodic boundary conditions, with the Hamiltonian given by
@@
Recall that since $A_{s}\rho=\rho A_{s}=\rho$ and $B_{p}\rho=\rho B_{p}=\rho$, ground states of Eq. (8) can be interpreted as closed loop condensates of the $e$, $m$, and $f$ anyons. After maximal decoherence, we instead only have $B_{p}\rho_{e}=\rho_{e}B_{p}=\rho_{e}$ and $A_{s}\rho_{e}A_{s}=\rho_{e}$. Physically, the $e$-noise has the effect of “freezing” the $m$-loops (and hence also the $f$ loops) into a classical ensemble, while leaving the “quantum” condensate of $e$-loops untouched. More precisely, $e$-errors break the strong 1-form magnetic symmetry of the original pure state down to a weak 1-form symmetry, while leaving the strong 1-form electric symmetry intact. We will later provide a general discussion of the role 1-form symmetries play in characterizing generic imTOs (see Sec. VI).
@@
Similar considerations hold for $\rho_{f}$. We have that $A_{s}B_{s-\mathbf{y}}\rho_{f}=\rho_{f}A_{s}B_{s-\mathbf{y}}=\rho_{f}$, where $A_{s}B_{s-\mathbf{y}}$ generates a closed $f$ loop and $s-\mathbf{y}$ denotes the plaquette to the south-east of vertex $s$, while we only have that $A_{s}\rho_{f}A_{s}=\rho_{f}$. Thus, $f$-noise freezes both the $e$ and $m$ loops but leaves the $f$ loops untouched, such that $\rho_{f}$ describes a quantum condensate of fermionic excitations (stated otherwise, $\rho_{f}$ retains a strong 1-form symmetry): this is not expected to occur in the ground state of a gapped, local 2D Hamiltonian. Remarkably, despite starting with a bosonic topological order, decoherence has resulted in a state effectively described as a condensate of fermions. In a rough sense, local decoherence allows one to “peel off” half of the original state.
@@
which yields the stabilizer group $\mathcal{S}_{f}=\langle A_{s}B_{s-\mathbf{y}},W_{x,y}^{f}\rangle$. This precisely describes a topological subsystem code describing the $\mathbb{Z}_{2}^{(1)}$ topological order which, again, is non-modular [40]. Like the previous case, $f$ is transparent (it braids trivially with itself) and hence the decohered state encodes no quantum memory, consistent with Ref. [31].
@@
It is readily apparent that the resulting TO in either case is that of a single $\mathbb{Z}_{2}$ Toric code, with the fusion group of the Abelian anyons given by $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. This can also be directly verified with the explicit forms of |$\ket{\rho_{e,f}}\rangle$ in the lattice model. In light of our above stabilizer analysis however, we note a key distinction between the mixed-states $\mathcal{C}_{e}$ and $\mathcal{C}_{f}$. Restricting attention to anyons with support on only the ket or bra space, we see that both orders support a single such anyon. For $\mathcal{C}_{e}$, this is the boson $e_{+}\sim e_{-}$, while for $\mathcal{C}_{f}$, this is the fermion $f_{+}\sim f_{-}$, where the equivalences are up to fusion with the condensed anyon. This is consistent with our observation in the stabilizer analysis that under $e$ and $f$ noise, the sole remaining coherent closed loops are simply those corresponding to the original $e$ and $f$ anyons, respectively.
@@
We now show that this anyon condensation across the ket and bra spaces, at the level of the density matrix in the original Hilbert space, corresponds to gauging out an anyon. Recall that anyon condensation proceeds in two steps (in Abelian theories): to condense an anyon $a$, one first (i) projects out from the theory those anyons which braid non-trivially with $a$ (i.e. they become confined) and then (ii) identifies those anyon types which differ by fusion with $a$. For instance, in the Toric code, condensing $e$ confines the $m$ and $f$ anyons while $e$ becomes identified with the vacuum: the resulting state has no remaining anyon excitations and is trivial. Gauging out an anyon however corresponds to only performing step (i) of this process. For instance, gauging out $e$ still confines $m$ and $f$, but leaves $e$ distinct from the vacuum, such that one is left with the anyon content $\{1,e\}$—precisely that of the $\mathbb{Z}_{2}^{(0)}$ non-modular TO realized via decoherence of $e$.
@@
Crucially, one can also gauge out anyons that cannot be condensed (i.e. non-bosonic anyons); analogously, one can decohere non-bosonic anyons $a$ since this corresponds to the conventional condensation of the bosonic pair $(a_{+}a_{-})$ in the doubled Hilbert space. For instance, one may gauge out $f$ from the $\mathbb{Z}_{2}$ Toric code to obtain the $\mathbb{Z}_{2}^{(1)}$ TO, precisely replicating the effect of $f$ errors. Surprisingly, as this simple example illustrates, locally correlated errors (which correspond to anyon condensation under the Choi map) provide a physical implementation of the gauging out procedure, which thus far remained a conceptual device for generating topological subsystem codes from topological stabilizer codes [32, 33, 34, 35, 40].
@@
Let us pause to recapitulate our observations in the context of the Toric code. We found that anyonic decoherence, previously shown to correspond to anyon condensation across the ket and bra spaces, implements the gauging out of anyons in the original Hilbert space, including those which are forbidden from condensing under purely unitary evolution. This process led to mixed-states supporting non-modular topological order (corresponding to topological subsystem codes) which is believed to be forbidden in the ground state of a locally gapped Hamiltonian in 2D. Thus, we claim that both the $e$- and $f$-decohered Toric codes represent intrinsically mixed topological states of matter. However, there is an important distinction between the two cases: namely, the strong 1-form symmetry of the $e$-decohered Toric code is non-anomalous, while that of the $f$-decohered Toric code is anomalous. In Sec. VI, we will discuss the implications of these anomalies and the sense in which we expect them to characterize mixed-state phases of matter. Note that for the specific case of the $\mathbb{Z}_{2}$ Toric code subject to $f$ errors, Ref. [31] numerically verified the robustness of the resulting imTO against finite noise channels that explicitly break the strong 1-form symmetry.
@@ III.2 $\mathbb{Z}_{4}^{(1)}$ TSSC from $\mathbb{Z}_{4}$ Toric code
$H_{\mathbb{Z}_{4}}=-\sum_{v}(A_{v}+A_{v}^{\dagger})-\sum_{p}(B_{p}+B_{p}^{\dagger}),$ (15)
with the star and plaquette operators defined in Fig. 2. The ground state manifold is determined by the constraints $A_{v}=B_{p}=1$, violations of which indicate the presence of electric $e$ and magnetic $m$ excitations, respectively. Explicitly, $Z_{i}$ applied to the ground state excites an $e$ and an $e^{-1}$ anyon at vertices connected by the edge $i$. Likwise, applying $X_{i}$ creates an $m$ and $m^{-1}$ on plaquettes separated by the edge $i$. These anyons satisfy $\mathbb{Z}_{4}\times\mathbb{Z}_{4}$ fusion rules $e^{4}=m^{4}=1$ and the braiding statistics between two composite objects $e^{a}m^{b}$ and $e^{c}m^{d}$ is given by $\mathcal{B}_{\theta}(ab,cd)=i^{ad+bc}$. On the torus, the non-contractible Wilson loops $W^{e}_{x,y}=\prod_{i\in\Gamma_{x,y}}Z_{i}$ and $W^{m}_{x,y}=\prod_{i\in\hat{\Gamma}_{x,y}}X_{i}$ serve as the logical operators and satisfy the algebra $W_{x/y}^{e}W_{y/x}^{m}=iW_{y/x}^{m}W_{x/y}^{e}$, such that the code space stores two $d=4$ qudits.
@@
Said more formally, in the language of higher-form symmetries to be discussed in Section VI, $\hat{\mathcal{A}}$ errors break most of the strong 1-form symmetries of $\rho$ down to weak 1-form symmetries, while leaving the strong 1-form symmetries corresponding to $\mathcal{A}$ anyons intact. This is encoded in Eq. (6) and the fact that the stabilizer group $\mathcal{S}_{e^{-1}m}$ for the TSSC is generated by small loops for the $em$ anyon (which generates $\mathcal{A}$). Finally, since $e^{2}m^{2}$ is transparent in $\hat{\mathcal{A}}$, it remains transparent in $\mathcal{A}$ by definition. We thus obtain the same result as above: the set of decohered density matrices on the torus form the code space for a TSSC which describes a non-modular Abelian anyon theory $\mathcal{A}$, whose non-trivial anyons are two semions and a transparent boson. The mutual statistics of the semions result in this mixed-state encoding a logical qubit in its logical subsystem. Since this mixed-state encodes non-trivial logical information, we expect on physical grounds that it is robust (up to a finite noise threshold) against finite-depth local quantum channels [57, 21] and so represents a genuine imTO phase of matter although, as emphasized above and in Section VI, this requires careful verification.
It is instructive to once again consider the gauging out procedure from the perspective of the doubled Hilbert space. In this picture, decoherence of the anyons in $\hat{\mathcal{A}}$ corresponds to condensing $\{[1]_{+}[1]_{-},$ $[e^{-1}m]_{+}[e^{-1}m]_{-},[e^{2}m^{2}]_{+}[e^{2}m^{2}]_{-},[em^{3}]_{+}[em^{3}]_{-}\}$, which form a Lagrangian subgroup of the TO in the doubled space $\mathcal{C}\times\bar{\mathcal{C}}$, where $\mathcal{C}=\mathbb{Z}_{4}\times\mathbb{Z}_{4}$. Anyon condensation proceeds in the usual way: each anyon from $\hat{\mathcal{A}}$ is identified with the vacuum. Next, any excitation which braids non-trivially with any condensed anyon becomes confined and, of the remaining deconfined excitations, any that differ only up to fusion by anyons in $\hat{\mathcal{A}}$ are identified. A simple calculation shows that the resulting topological order is that of a $\mathbb{Z}_{4}$ gauge theory, with only the following anyons supported solely on the ket space: $[em]_{+}$, and $[e^{-1}m^{-1}]_{+}$, while $[e^{2}m^{2}]_{+}\sim[e^{2}m^{2}]_{-}$ can move freely between the ket and bra spaces, and is a transparent anyon. As expected, these correspond precisely to those anyons in $\mathcal{A}$, obtained by applying the gauging out procedure in the original Hilbert space.
@@ III.3 Chiral Semion from Double Semion
As an instance of this mapping, let us take the doubled semion anyon theory as our parent TO. This theory can be realized as a Pauli stabilizer Hamiltonian [36] and its anyons form a $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ group under fusion, with elements $\{1,s,\bar{s},s\bar{s}\}$. Here, $s$ is a semion (it has self-statistics $\theta(s)=i$), $\bar{s}$ is an anti-semion ($\theta(\bar{s})=-i$), and $s\bar{s}$ is a boson. Now, we subject a ground state of this system to an error channel that incoherently proliferates the semion $s$ i.e., $\hat{\mathcal{A}}=\{1,s\}$. This corresponds to gauging out $s$, which braids trivially with $\bar{s}$. The resulting anyon theory for the decohered mixed-state is given by $\mathcal{A}=\{1,\bar{s}\}$ i.e., it is the chiral (anti)-semion Abelian anyon theory. Since $\bar{s}$ has non-trivial self-statistics, the decohered code space encodes exactly one logical qubit; this example represents the minimal model in which one obtains a chiral anyon theory with an encoded logical qubit. Again, the presence of a logical qubit confers stability to this state against finite-depth local channels and it represents a mixed-state phase of matter. While this mixed state does not correspond to an imTO, as it is described by a modular anyon theory, chiral UMTCs cannot arise in the ground states of locally commuting Hamiltonians [50] and thus cannot be realized in fixed point wave-functions (with finite-dimensional local Hilbert spaces). Nonetheless, here we have shown that a chiral UMTC can, in fact, arise in a mixed state.